OpenStudy (anonymous):

Use the point-slope form of the linear equation to find an equation of the line with the given slope and passing through the given point. Then write the equation in standard form. Slope -4; through (-7, -5)

OpenStudy (anonymous):

y=mx=b, now plug in the values -5=-4(-7)+b b=23 y=-4x+23

OpenStudy (anonymous):

a. 4x + y = -5 b. 4x + y = -27 c. 4x + y = -33 d. 4x + y = 33

OpenStudy (anonymous):

hmmm, I must have made a mistake somewhere, but I do not see it. Because mine in standard is 4x+y=23...did you copy the equation right?

OpenStudy (anonymous):

yup i copied it right!